I borrowed 1000 today if i need to pay annually for 10 years how much will my payment be if interest is : 15% compounded bi-monthly during the 1st five ears and 12% compounded continuously during the next five years?


1
Expert's answer
2021-10-12T12:27:08-0400
"\frac{1}{10}\times(1000(1+\frac{0.15}{24})^{10}+1000\times e^{0.6})=288.64"

"\"